A merchant has coffe worth $30 a pound that she wishes to mix with 40 pounds of coffee worth $90 a pound to get a mixture that can be sold for $70 a pound. How many pounds of the $30 coffee should be used?
Answer by mananth(16946)

-------- price ---------------- quantity
Coffee type I 90 ----------------40 pounds
Coffee type II 30 ----------------x pounds
Mixture 70 ----------------40+x pounds

90*40+30x=70(40+x)
3600+30x=2800+70x
30x-70x=2800-3600
-40x=-800
/-40
x=20pounds of Coffee type II

m.ananth@hotmail.ca